What is the theoretical yield of dehydroascorbic acid for the reaction of 1000.0 mg of vitamin C with 1.000 g of oxygen
1
Expert's answer
2020-05-05T03:40:05-0400

The oxidation of ascorbic acid can be described as following:

Ascorbic acid + 1/2 O2 --> Dehydroascorbic acid + H2O

The limiting reactant must be determined to find the he theoretical yield of dehydroascorbic acid. For teh purpose, the number of moles of reactants must be calculated.

According to the reaction:

n(ascorbic acid) = m / Mr = 1000.0 mg / 176 g/mol = 1 g / 176g = 0.006 mol

n(O2) = m / (1/2 × Mr) = 1.000 g / (1/2 × 32 g/mol) = 0.0625 mol

As n(ascorbic acid) < n(O2), ascorbic acid is a limiting reactant and must be used for the calculation of the mass of dehydroascorbic acid. As a result:

m(dehydroascorbic acid) = m(ascorbic acid) × Mr(dehydroascorbic acid) / Mr(ascorbic acid) = 1000.0 mg × 174 g/mol / 176 g/mol = 988.64 g


Answer: 988.64 g